OpenStudy (anonymous):

can someone help me with this question? Iris is cutting fleece to make scarves for her friends. She has thirteen and one-third feet of fleece to make 5 scarves. How long can she make each scarf? Show your work, and include cancellation in your steps

OpenStudy (lgbasallote):

dont you just divide?

OpenStudy (anonymous):

13 1/3 = 40/3 40/3 *1/5 =

OpenStudy (lgbasallote):

you're not really a @completeidiot , are you?

OpenStudy (anonymous):

lol i think thats just a name... right?

OpenStudy (anonymous):

its so people dont expect much from me

OpenStudy (anonymous):

oh ok lol thx for the help